### v1

@@ 5.4 Limitations
The main limitation of our method lies in setting the value for $\rho$. Our results show that setting a conservative value is sufficient for practice to avoid model damage while achieving SOTA unlearning performance. We argue that this problem is less relevant for practice and mainly a byproduct of the benchmark design. An attacker in real-life would not gain anything from using a poison trigger that redirects to the original label. $\rho$ should therefore be trivial to set in practice with a small buffer for potential clean and poison label overlaps.
@@ 6 Conclusion
We present a novel method for poison trigger unlearning, [Uncaptioned image] XLF, that addresses the challenges of hyperparameter selection and model accuracy retention in the absence of ground truth. By effectively unlearning poison triggers while preserving model performance, even when only a subset of the poisoned data is identified, our approach significantly advances the state of the art in mitigating adversarial attacks on machine learning models. Our evaluation on standard benchmark datasets proposed by Goel et al. (2024) demonstrates the performance improvements of our method in both poison removal and model protection compared to existing methods such as SSD and full retraining from scratch. The proposed method holds promise for enhancing the robustness and resilience of machine learning systems deployed in real-world scenarios where adversarial attacks are a growing concern. Future work will focus on extending these methods to larger and more complex models and exploring their applicability to different types of poison attacks.

### v2

@@ 5.5 Limitations
A limitation of our method lies in setting the value for $\rho$. Our results show that setting a conservative value is sufficient for practice to avoid model damage while achieving SOTA unlearning performance. We argue that this problem is less relevant for practice and mainly a byproduct of the benchmark design. An attacker in real-life would not gain anything from using a poison trigger that redirects to the original label. $\rho$ should therefore be trivial to set in practice with a small buffer for potential clean and poison label overlaps. The experimental results on Imagenette and ILSVCR Imagenet using a ViT further show that our method works well without changing parameters across vastly different datasets, model architectures and poison types.
While we set a new SOTA, our experiments show two key challenges for future work to advance the field of poison unlearning:
  1. 1.
    Closing the gap to 100% poison healing: For safety in practice methods should be able to unlearn (near) 100% of poison, as any remaining high-severity poison trigger could cause significant damage.
  2. 2.
    Improved generalisation from limited samples: While our method successfully unlearns a majority of the poisoned data given only a subset of the poison, further generalisation is necessary to unlearn more complex poisons. To one-shot heal a moving poison trigger, an unlearning method needs to identify the trigger and then be able to correctly remove it from the model regardless of its location in the image. This is beyond the capability of current methods.
@@ 6 Conclusion
We present a novel method for poison trigger unlearning, [Uncaptioned image] XLF, that addresses the challenges of hyperparameter selection and model accuracy retention in the absence of ground truth. By effectively unlearning poison triggers while preserving model performance, even when only a subset of the poisoned data is identified, our approach significantly advances the state of the art in mitigating adversarial attacks on machine learning models. Our evaluation on standard benchmark datasets proposed by Goel et al. (2024) demonstrates the performance improvements of our method in both poison removal and model protection compared to existing methods such as SSD and full retraining from scratch. The proposed method holds promise for enhancing the robustness and resilience of machine learning systems deployed in real-world scenarios where adversarial attacks are a growing concern. Future work will focus on closing the gap to unlearning 100% of poison and improved generalisation capabilities of unlearning methods.
Broader Impact Statement
Our work adds positive impact by providing model owners with better methods to remove adversarial attacks from already trained models that could harm users. Given the increasing reliance on data from openly accessible and cheaply manipulable sources on the web, our contribution is expected to grow in importance. We see two main negative impacts of our work. First, a false sense of trust in our methods can lead model owners into a false sense of security, neglecting data pre-filtering as they can ”just fix it afterwards”, or failing to remove poison leading to harm. Second, by sharing our method openly accessible, attackers can devise new attacks that bypass our defences. We see these negative impacts in line with general AI and cyber-security. The first problem is addressed by clearly showing the imperfections and limitations of our work. The second problem is the eternal cat-and-mouse game in any security-related field. Overall, we see the positive impact of our research outweighing the risks.
Acknowledgments and Disclosure of Funding
No competing interests. This work was supported by the Accenture Turing Strategic Partnership, the Turing Enrichment scheme, EPSRC CDT AgriFoRwArdS [grant number EP/S023917/1], and EPSRC DTP [grant number EP/W524633/1].